# The Chicago-Carnegie Hubble Program:

### v2

@@ 3.5 Galaxies in which the JAGB magnitude failed to converge
We also emphasize that because we performed this analysis blinded, we discarded these three galaxies from our SN Ia calibrating sample before knowing their measured distances. Our blinding procedure is described in detail in the CCHP $H_{0}$ overview paper in Freedman et al. (2024).
@@ 3.5.4 Distances to NGC 3972, NGC 4038, NGC 4424
Although NGC 3972, NGC 4038, and NGC 4424 never technically converged according to our algorithm, we still visually selected ‘outer disk‘ radial cuts to provide distances to these galaxies. Then, we re-measured $H_{0}$ while including these galaxies in the SN Ia calibration. The distances to these three galaxies and their uncertainties are presented in Tables 3 and 4, although they were unused for our ‘primary’ Hubble constant measurement. As described in Section 3.8 and in Freedman et al. (2024), we re-measured $H_{0}$ by including NGC 4038 and NGC 3972 to quantify the effect of their exclusion in our primary Hubble constant measurement. NGC 4424, which has almost a nonexistent intermediate-age population and a JAGB distance that is clearly an outlier compared with the Cepheid and TRGB measured distances, was excluded entirely.
While the JAGB magnitude in NGC 3972 never converged, as shown in Figure 7, we selected a radial cut of 9.0 kpc. In NGC 4038, while the JAGB magnitude technically converged at $d\approx 20$ kpc, the presence of NGC 4039 from $3828<d<$ 38 kpc made this galaxy pair an extraordinary exception to our algorithm. Nevertheless, we selected a radial cut of 42.0 kpc. Finally, in NGC 4424, we selected a radial cut of 7.3 kpc. While the JAGB magnitude indeed converged in the outer disk, only a few hundred JAGB stars remained so we completely rejected this galaxy from both our primary and secondary SN Ia galaxy samples.
@@
M101 29.22 0.03 0.03 7.0 1279
NGC 1365 31.39 0.02 0.03 19.0 1242
NGC 2442 31.61 0.01 0.04 21.0 931
NGC 4536 30.98 0.01 0.03 15.7 2824
NGC 4639 31.74 0.02 0.03 22.1 1292
NGC 5643 30.59 0.01 0.04 13.1 1098
NGC 7250 31.60 0.02 0.04 20.9 848
NGC 3972 31.68 0.04 0.03 21.7 829
NGC 4038 31.54 0.06 0.04 20.3 435
NGC 4424 31.16 0.04 0.03 17.1 381
@@ 3.6.1 Statistical Uncertainties
    We note the mode measured from an increasingly smoothed luminosity function will eventually converge to the mean. Therefore, this ‘smoothing parameter uncertainty’ encapsulates any systematic differences incurred from the choice of JAGB statistic (i.e., mode vs. mean vs. median). We demonstrate in Appendix B that if we use the mean instead of the mode as the chosen JAGB statistic, the distance moduli were measured to be $0.035$ mag brighter on average. This systematic offset is fully encapsulated within the smoothing parameter uncertainty of our zeropoint NGC 4258 (0.05 mag).
@@ 3.7 Adopted Zero-point Calibration in NGC 4258
Choice of $\sigma_{s}$  $\cdots$ 0.05  $\cdots$
@@
The JAGB magnitude in the combined inner and outer disk JAGB sample was measured to be $m_{JAGB}=23.43$ mag. We corrected this for a Milky Way foreground extinction of $A_{F115W}=0.02$ mag (Schlegel et al., 1998). This yielded a JAGB zeropoint of $M_{JAGB}=-5.99\pm 0.05$ (stat) ±$\pm~{}0.04$ (sys) mag. The sources of uncertainty on this zeropoint are summarized in Table 5. In addition, we adopted a 1% systematic error (0.02 mag) for the uncertainty on the NIRCam photometric zero-point. This uncertainty came from STScI, which currently quotes the NIRCam absolute flux uncertainties at $\lesssim 1\%$. [^9] https://jwst-docs.stsci.edu/jwst-calibration-status/nircam-calibration-status/nircam-imaging-calibration-status
@@ 3.8 Measurement of the Hubble constant
Details of the JAGB calibration of SNe Ia, as well as comparisons between the JAGB distances measured in this paper with TRGB and Cepheid distances, can be found in the CCHP $H_{0}$ overview paper Freedman et al. (2024). We give a brief summary here.
Freedman et al. (2024) applied the JAGB distances measured in this paper to two samples of distant supernovae: our primary SNe Ia sample, the Carnegie Supernova Program (CSP; Uddin et al. 2024), and the SNe Ia sample used by the SHoES team, (2) the Pantheon+ sample (Scolnic et al., 2022). In our sample of seven galaxies, there are eight SN Ia calibrators that have been observed as part of Pantheon+ and seven for the CSP (note there are two SNe Ia observed in NGC 5643 but one of them, SN 2017cbv, was discovered after the CSP program ended). These distances were used as inputs to the MCMC analysis described in Freedman et al. (2024) for the CSP sample. For the Pantheon+ analysis, we determined $H_{0}$ by scaling the $M_{B}$ values determined by the JAGB method relative to those measured by SHoES, instead of completely reanalyzing the entire SN Ia sample again. Thus, the $H_{0}$ determined from Pantheon+ was measured as a cross-check, but lacks measured uncertainties.
Freedman et al. (2024) found excellent agreement between the CSP and Pantheon+ determined Hubble constants. The CSP JAGB $H_{0}$ (our primary measurement) was determined to be $H_{0}=67.96\pm 1.85~{}\rm{(stat)}\pm 1.90~{}\rm{(sys)}~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$. The Pantheon+ JAGB $H_{0}$ was determined to be $H_{0}=68.56~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$. Therefore, the differing SN Ia samples alone cannot fully explain the Hubble tension.
In Freedman et al. (2024), we also measured a Hubble constant by including NGC 3972 and NGC 4038 as calibrators, as discussed in Section 3.5.4. Including these two galaxies delivered a Hubble constant of $H_{0}=68.76\pm 1.86~{}\rm{(stat)\pm 1.90~{}\rm{(sys)}~{}km~{}s^{-1}~{}Mpc^{-1}}$.
@@ 4 Summary & Future Outlook
We used JWST imaging to measure JAGB distances to seven SN Ia host galaxies, performing our analysis completely blinded. We determined a value of the Hubble constant of $H_{0}=67.96\pm 1.85~{}\rm{(stat)}\pm 1.90~{}\rm{(sys)}~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$ (Freedman et al., 2024). This value is in excellent agreement with the CMB value measured by Planck of $H_{0}=67.4\pm 0.5~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$ (Planck Collaboration et al., 2020). The JAGB distances measured in this paper can cross-check the distances to the same galaxies measured via Cepheids and the TRGB (using the same JWST imaging), respectively presented in Owens et al. (2024); Hoyt et al. (2024), both in preparation. We present these comparisons in our CCHP JWST $H_{0}$ overview paper, Freedman et al. (2024). In summary, the CCHP TRGB and JAGB distances agreed well, with the weighted mean difference between the JAGB minus TRGB distance moduli being $+0.017\pm 0.031$ mag. On the other hand, our JAGB distance moduli disagreed at the $3\sigma$ level with the Cepheid distances, where the weighted mean difference between the JAGB - Cepheid distance moduli was measured to be $+0.086\pm 0.028$ mag. Therefore, a significant result from this JWST program is that the two distance indicators least impacted by crowding/blending (TRGB and JAGB) agree well with each other, but both disagree with the Cepheid distance indicator, which uses stars in the crowded, high surface-brightness disks of galaxies and consistently measures distances to be closer. Understanding the reasons for this ‘local Hubble tension’ disagreement will be essential before confirming the larger Hubble tension between the local and early Universe measurements of $H_{0}$.
@@
    JWST. The JWST will continue to provide imaging of SN Ia host galaxies. We will continue to measure their JAGB distances to increase the number of Type Ia supernovae calibrated by the JAGB method. In particular, a major conclusion from Freedman et al. (2024) was that the SN Ia absolute magnitudes calibrated by Cepheids determined in Riess et al. (2022) grow fainter with increasing distance, an effect potentially resulting from increasing crowding/blending effects. Therefore, additional JWST data of farther SN Ia host galaxies, particularly at distances beyond 40 Mpc, will be imperative in definitively ruling out systematic effects in the local distance scale measurement of $H_{0}$.
@@
Appendix A ‘Inner Disk’ Color-magnitude diagrams
Appendix B Mean vs. Median vs. Mode
Adopting the mean resulted in distance moduli that were $0.035\pm 0.010$ mag closer than distance moduli measured using the mode. This corresponds to a 1.6% larger $H_{0}$. Adopting the median resulted in distance moduli that were $0.033\pm 0.007$ mag closer than distance moduli measured using the mode. This corresponds to a 1.5% larger $H_{0}$. These systematic differences were fully encapsulated within our minimum adopted smoothing parameter error of $0.05$ mag,101010 The total smoothing parameter error, which adds in quadrature the smoothing parameter error from the SN Ia host galaxy and NGC 4258, ranged from 0.05 to 0.06 mag for all the host galaxies. because measuring the mode of an increasingly smoothed distribution will eventually converge to measuring the mean of that distribution, as discussed in Section 3.6.1. Therefore, using the mode as the JAGB statistic is equivalent to using the mean/median as long as an appropriate smoothing parameter error is adopted. In conclusion, the differing shapes of the JAGB LFs in our sample propagate to a $\sim 1.5$% error to the final measured $H_{0}$, which we accounted for by adopting a minimum 2% uncertainty in our error budget.
Appendix C Masking spiral arms
@@
If we chose to keep the spiral arms unmasked, the JAGB magnitudes in M101, NGC 2442, NGC 4258, and NGC 4639 would have been instead respectively measured to be: 23.21 (0.03 mag brighter), 25.79 (0.01 mag brighter), 23.43 (same magnitude), and 25.76 (0.01 mag brighter). This resulted in distance moduli that were on average 0.007 mag brighter for the seven SN Ia host galaxies, or a 0.3% larger $H_{0}$.

### v3

@@ 3.5 Galaxies in which the JAGB magnitude failed to converge
We also emphasize that because we performed this analysis blinded, we discarded these three galaxies from our SN Ia calibrating sample before knowing their measured distances. It was at this point, after finalizing the analysis and SN Ia calibrating sample, that we removed the random photometric offset from our catalogs to reveal each galaxy’s measured JAGB distance.
@@ 3.5.4 Distances to NGC 3972, NGC 4038, NGC 4424
Although NGC 3972, NGC 4038, and NGC 4424 never technically converged according to our algorithm, we still visually selected ‘outer disk‘ radial cuts to provide distances to these galaxies. Then, we re-measured $H_{0}$ while including these galaxies in the SN Ia calibration. The distances to these three galaxies and their uncertainties are presented in Tables 3 and 4, although they were unused for our ‘primary’ Hubble constant measurement. As described in Section 3.8 and in Freedman et al. (2025), we re-measured $H_{0}$ by including NGC 4038, NGC 3972, and NGC 4424 to quantify the effect of their exclusion in our primary Hubble constant measurement.
While the JAGB magnitude in NGC 3972 never converged, as shown in Figure 7, we selected a radial cut of 9.0 kpc. In NGC 4038, while the JAGB magnitude technically converged at $d\approx 20$ kpc, the presence of NGC 4039 from $3828<d<$ 38 kpc made this galaxy pair an extraordinary exception to our algorithm. Nevertheless, we selected a radial cut of 42.0 kpc. Finally, in NGC 4424, we selected a radial cut of 7.3 kpc.
@@
M101 29.21 0.03 0.03 7.0 1279
NGC 1365 31.38 0.02 0.03 19.0 1242
NGC 2442 31.60 0.01 0.04 21.0 931
NGC 4536 30.97 0.01 0.03 15.7 2824
NGC 4639 31.73 0.02 0.03 22.1 1292
NGC 5643 30.58 0.01 0.04 13.1 1098
NGC 7250 31.59 0.02 0.04 20.9 848
NGC 3972 31.67 0.04 0.03 21.7 829
NGC 4038 31.53 0.06 0.04 20.3 435
NGC 4424 31.15 0.04 0.03 17.1 381
Note. — Note these distance uncertainties do not include the uncertainties from the JAGB zero-point listed in Table 5.
@@ 3.6.1 Statistical Uncertainties
    This uncertainty calculation assumes the distribution of JAGB stars is Gaussian. To test whether this is a valid assumption for calculating the statistical uncertainty, we sub-divided the final distribution of $\sim 2800$ JAGB stars in NGC 4536, the galaxy with the largest number of JAGB stars, into five radially-separated bins, each with $\sim 570$ JAGB stars. Because the error on the mode for this total sample was measured to be 0.006 mag, the expected uncertainty in each sub-region should be $\sqrt{n/k}\times 0.006$, where $n$ is 2800 and $k$ is the number of JAGB stars in each subregion. We then measured the error on the mode in each region, finding: 0.015, 0.014, 0.014, 0.013, 0.013 mag in each subregion, compared to the expected error of 0.014 mag. This test validated the scatter is close to Gaussian and that our method for using the error on the mode as a statistical uncertainty is applicable and valid.
@@
    We note the mode measured from an increasingly smoothed luminosity function will eventually converge to the mean. Therefore, this ‘smoothing parameter uncertainty’ encapsulates any systematic differences incurred from the choice of JAGB statistic (i.e., mode vs. mean vs. median). We demonstrate in Appendix C that if we use the mean instead of the mode as the chosen JAGB statistic, the distance moduli were measured to be $0.030$ mag brighter on average. This systematic offset is fully encapsulated within the smoothing parameter uncertainty of our zeropoint NGC 4258 (0.04 mag).
@@ 3.7 Adopted Zero-point Calibration in NGC 4258
Choice of $\sigma_{s}$  $\cdots$ 0.04  $\cdots$
@@
The JAGB magnitude in the combined inner and outer disk JAGB sample was measured to be $m_{JAGB}=23.44$ mag. We corrected this for a Milky Way foreground extinction of $A_{F115W}=0.02$ mag (Schlegel et al., 1998). This yielded a JAGB zeropoint of $M_{JAGB}=-5.98\pm 0.05$ (stat) ±$\pm~{}0.04$ (sys) mag. The sources of uncertainty on this zeropoint are summarized in Table 5. In addition, we adopted a 1% systematic error (0.02 mag) for the uncertainty on the NIRCam photometric zero-point. This uncertainty came from STScI, which currently quotes the NIRCam absolute flux uncertainties at $\lesssim 1\%$. [^10] https://jwst-docs.stsci.edu/jwst-calibration-status/nircam-calibration-status/nircam-imaging-calibration-status
We also measured the JAGB magnitude in the inner and outer fields separately, after adopting the radial cut of $d>7.3$ kpc. In the inner field, we measured a JAGB magnitude of $m_{JAGB}=23.41\pm 0.04$ (stat) mag, with $\sim 2300$ JAGB stars contributing to the measurement. In the outer field, we measured a JAGB magnitude of $m_{JAGB}=23.51\pm 0.06$ (stat) mag, with $\sim 600$ JAGB stars contributing to the measurement. The statistical uncertainty was calculated by adding in quadrature the the error on the mode and the smoothing parameter uncertainty. These two measurements agree to within $1.2\sigma$.
@@ 3.8 Measurement of the Hubble constant
Details of the JAGB calibration of SNe Ia, as well as comparisons between the JAGB distances measured in this paper with TRGB distances, can be found in the CCHP $H_{0}$ results paper Freedman et al. (2025). We give a brief summary here.
Freedman et al. (2025) applied the JAGB distances measured in this paper to SNe Ia in the Carnegie Supernova Program (CSP; Uddin et al. 2024). In our sample of seven galaxies, there are eight SN Ia calibrators (NGC 5643 contains two SN Ia). These distances were used as inputs to the MCMC analysis described in Freedman et al. (2025) for the CSP sample. The CSP JAGB $H_{0}$ was determined to be $H_{0}=67.80\pm 2.17~{}\rm{(stat)}\pm 1.64~{}\rm{(sys)}~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$.
In Freedman et al. (2025), we also measured a Hubble constant by including NGC 3972, NGC 4038, and NGC 4424 as calibrators, as discussed in Section 3.5.4. Including these three additional SN Ia delivered a Hubble constant of $H_{0}=68.75\pm 2.23~{}\rm{(stat)}\pm 1.65~{}\rm{(sys)}~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$.
@@ 4 Summary & Future Outlook
We used JWST imaging to measure JAGB distances to seven SN Ia host galaxies, performing our analysis completely blinded. We determined a value of the Hubble constant of $H_{0}=67.80\pm 2.17~{}\rm{(stat)}\pm 1.64~{}\rm{(sys)}~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$ (Freedman et al., 2025). This value is in excellent agreement with the CMB value measured by Planck of $H_{0}=67.4\pm 0.5~{}\rm{km~{}s^{-1}~{}Mpc^{-1}}$ (Planck Collaboration et al., 2020). The JAGB distances measured in this paper can cross-check the distances to the same galaxies measured via Cepheids and the TRGB (using the same JWST imaging), respectively presented in Hoyt et al. (2025) and Owens et al. (in preparation). We present these comparisons in our CCHP JWST $H_{0}$ results paper, Freedman et al. (2025).
@@
    JWST. The JWST will continue to provide imaging of SN Ia host galaxies. We will continue to measure their JAGB distances to increase the number of Type Ia supernovae calibrated by the JAGB method. As discussed in Freedman et al. (2025), more JAGB distances to SN hosts are needed to provide a more representative sample of SNe Ia. Therefore, additional JWST data of farther SN Ia host galaxies, particularly at distances beyond 40 Mpc, will be imperative in definitively ruling out systematic effects in the local distance scale measurement of $H_{0}$.
@@
Appendix A Masking spiral arms
@@
If we chose to keep the spiral arms unmasked, the JAGB magnitudes in M101, NGC 2442, NGC 4258, and NGC 4639 would have been instead respectively measured to be: 23.21 (0.03 mag brighter), 25.79 (0.01 mag brighter), 23.44 (same magnitude), and 25.76 (0.01 mag brighter). This resulted in distance moduli that were on average 0.007 mag brighter for the seven SN Ia host galaxies, or a 0.3% larger $H_{0}$.
Appendix B ‘Inner Disk’ Color-magnitude diagrams
Appendix C Mean vs. Median vs. Mode
Adopting the mean resulted in distance moduli that were $0.030\pm 0.010$ mag closer than distance moduli measured using the mode. This corresponds to a 1.4% larger $H_{0}$. Adopting the median resulted in distance moduli that were $0.026\pm 0.007$ mag closer than distance moduli measured using the mode. This corresponds to a 1.2% larger $H_{0}$. These systematic differences were fully encapsulated within our minimum adopted smoothing parameter error of $0.04$ mag because measuring the mode of an increasingly smoothed distribution will eventually converge to measuring the mean of that distribution, as discussed in Section 3.6.1. Therefore, using the mode as the JAGB statistic is equivalent to using the mean/median as long as an appropriate smoothing parameter error is adopted. In conclusion, the differing shapes of the JAGB LFs in our sample propagate to a $\sim 1.4$% error to the final measured $H_{0}$, which we accounted for by adopting a minimum 2% uncertainty in our error budget.